# On the Lipschitz Continuity of Set Aggregation Functions and Neural Networks for Sets

### v2

@@ A.4.3 The max function is Lipschitz continuous with respect to the matching distance
Note that $\forall k\in[d]$, there exists at least one $\mathbf{v}_{i}\in X$ such that $[\mathbf{v}_{\text{max}}]_{k}=[\mathbf{v}_{i}]_{k}$, and also at least one $\mathbf{u}_{j}\in Y$ such that $[\mathbf{u}_{\text{max}}]_{k}=[\mathbf{u}_{j}]_{k}$. Furthermore, we denote by $\pi^{*}$ the matching produced by the solution of the matching distance function:
$\pi^{*}=\operatorname*{arg\,min}_{\pi\in\mathfrak{S}_{M}}\sum_{i=1}^{M}||\mathbf{v}_{\pi(i)}-\mathbf{u}_{i}||$
Then, $\forall k\in[M]$ where $[\mathbf{v}_{\text{max}}]_{k}\geq[\mathbf{u}_{\text{max}}]_{k}$ and $[\mathbf{v}_{\text{max}}]_{k}=[\mathbf{v}_{\pi(i)}]_{k}$, we have that:
$\big{(}[\mathbf{v}_{\text{max}}]_{k}-[\mathbf{u}_{\text{max}}]_{k}\big{)}^{2}$ $=\big{(}[\mathbf{v}_{\pi(i)}]_{k}-[\mathbf{u}_{\text{max}}]_{k}\big{)}^{2}$
$=\min_{j\in[M]}\big{(}[\mathbf{v}_{\pi(i)}]_{k}-[\mathbf{u}_{j}]_{k}\big{)}^{2}$ (3)
$\leq\big{(}[\mathbf{v}_{\pi(i)}]_{k}-[\mathbf{u}_{i}]_{k}\big{)}^{2}$
Likewise, $\forall k\in[M]$ where $[\mathbf{u}_{\text{max}}]_{k}>[\mathbf{v}_{\text{max}}]_{k}$ and $[\mathbf{u}_{\text{max}}]_{k}=[\mathbf{u}_{i}]_{k}$, we have that:
@@
$\leq\big{(}[\mathbf{v}_{\pi(i)}]_{k}-[\mathbf{u}_{i}]_{k}\big{)}^{2}$
Then, from equations (2), (A.4.3), (A.4.3), and assuming that $[\mathbf{u}_{\text{max}}]_{1}=[\mathbf{u}_{1}]_{1}\geq[\mathbf{v}_{\text{max}}]_{1}$, $[\mathbf{u}_{\text{max}}]_{2}=[\mathbf{u}_{1}]_{2}\geq[\mathbf{v}_{\text{max}}]_{2}$ and $[\mathbf{v}_{\text{max}}]_{d}=[\mathbf{v}_{\pi(M)}]_{d}\geq[\mathbf{u}_{\text{max}}]_{d}$, we have that:
$\Big{|}\Big{|}f_{\text{max}}(X)-f_{\text{max}}(Y)\Big{|}\Big{|}$ $=\sqrt{\big{(}[\mathbf{v}_{\text{max}}]_{1}-[\mathbf{u}_{\text{max}}]_{1}\big{)}^{2}+\big{(}[\mathbf{v}_{\text{max}}]_{2}-[\mathbf{u}_{\text{max}}]_{2}\big{)}^{2}+\ldots+\big{(}[\mathbf{v}_{\text{max}}]_{d}-[\mathbf{u}_{\text{max}}]_{d}\big{)}^{2}}$
$\leq\sqrt{\big{(}[\mathbf{v}_{\pi(1)}]_{1}-[\mathbf{u}_{1}]_{1}\big{)}^{2}+\big{(}[\mathbf{v}_{\pi(1)}]_{2}-[\mathbf{u}_{1}]_{2}\big{)}^{2}+\ldots+\big{(}[\mathbf{v}_{\pi(M)}]_{d}-[\mathbf{u}_{M}]_{d}\big{)}^{2}}$
$\leq\sqrt{\big{(}[\mathbf{v}_{\pi(1)}]_{1}-[\mathbf{u}_{1}]_{1}\big{)}^{2}+\big{(}[\mathbf{u}_{\pi(2)}]_{2}-[\mathbf{v}_{2}]_{2}\big{)}^{2}}+\ldots+\sqrt{\big{(}[\mathbf{v}_{\pi(M)}]_{d}-[\mathbf{u}_{M}]_{d}\big{)}^{2}}$
$\leq||\mathbf{v}_{\pi(1)}-\mathbf{u}_{1}||+||\mathbf{v}_{\pi(2)}-\mathbf{u}_{2}||+\ldots+||\mathbf{v}_{\pi(M)}-\mathbf{u}_{M}||$
@@
A.4.4 The max function is Lipschitz continuous with respect to EMD
Let $\mathcal{X}$ denote a set that contains multisets of vectors of equal cardinalities, i. e., $|X|=M$, $\forall X\in\mathcal{X}$ where $M\in\mathbb{N}$. Let $X,Y\in\mathcal{X}$ denote two multisets. We have shown in subsection A.4.3 above that:
$\Big{|}\Big{|}f_{\text{max}}(X)-f_{\text{max}}(Y)\Big{|}\Big{|}\leq d_{M}(X,Y)$
@@ A.4.4 The max function is Lipschitz continuous with respect to EMD
A.5 Proof of Theorem 3.3
A.5.1 There exist $\textsc{NN}_{\textsc{sum}}$ models which are not Lipschitz continuous with respect to the matching distance
$\mathbf{v}_{X}=f_{2}\bigg{(}\text{ReLU}\Big{(}\big{\{}\!\!\big{\{}f_{1}(\mathbf{v}_{1}),\ldots,f_{1}(\mathbf{v}_{m})\big{\}}\!\!\big{\}}\Big{)}\bigg{)}$
where $X=\{\!\!\{\mathbf{v}_{1},\ldots,\mathbf{v}_{m}\}\!\!\}$ is a multiset, and $f_{1}$ and $f_{2}$ are fully-connected layers, i. e., $f_{1}(x)=a_{1}\,x+b_{1}$ and $f_{2}(x)=a_{2}\,x+b_{2}$. Furthermore, suppose that $a_{1}>0$, $b_{1}>0$ and $a_{2}>0$. Note that the Lipschitz constants of $f_{1}$ and $f_{2}$ are $\text{Lip}(f_{1})=a_{1}$ and $\text{Lip}(f_{2})=a_{2}$, respectively.
$\bigg{|}\bigg{|}f_{2}\bigg{(}\sum_{j=1}^{2}f_{1}(c)\bigg{)}-f_{2}\bigg{(}f_{1}(c)\bigg{)}\bigg{|}\bigg{|}$ $=\bigg{|}\bigg{|}f_{2}\bigg{(}\sum_{j=1}^{2}\text{ReLU}(a_{1}\,c+b_{1})\bigg{)}-f_{2}\bigg{(}\text{ReLU}(a_{1}\,c+b_{1})\bigg{)}\bigg{|}\bigg{|}$
$=\Big{|}\Big{|}f_{2}\Big{(}2\,(a_{1}\,c+b_{1})\Big{)}-f_{2}\Big{(}a_{1}\,c+b_{1}\Big{)}\Big{|}\Big{|}$
$=\Big{|}\Big{|}a_{2}\Big{(}2\,(a_{1}\,c+b_{1})\Big{)}+b_{2}-a_{2}\Big{(}a_{1}\,c+b_{1}\Big{)}-b_{2}\Big{|}\Big{|}$
$=||a_{2}(a_{1}\,c+b_{1})||$
$>||a_{2}(a_{1}\,c+c\,L\,a_{1})||$
$=||(L+1)\,a_{2}\,a_{1}\,c||$
$>(L+1)\,\text{Lip}(f_{2})\,\text{Lip}(f_{1})\,c$
$>L\,\text{Lip}(f_{2})\,\text{Lip}(f_{1})\,d_{M}(X,Y)$

### v3

@@ B.4.5 The max function is Lipschitz continuous with respect to the matching distance
Note that $\forall k\in[d]$, there exists at least one $\mathbf{v}_{i}\in X$ such that $\mathbf{v}_{\text{max}}]_{k}=[\mathbf{v}_{i}]_{k}$, and also at least one $\mathbf{u}_{j}\in Y$ such that $\mathbf{u}_{\text{max}}]_{k}=[\mathbf{u}_{j}]_{k}$. Furthermore, we denote by $\pi^{*}$ the matching produced by the solution of the matching distance function:
$\pi^{*}=\operatorname*{arg\,min}_{\pi\in\mathfrak{S}_{M}}\sum_{i=1}^{M}\|\mathbf{v}_{\pi(i)}-\mathbf{u}_{i}\|$
Then, $\forall k\in[d]$ where $\mathbf{v}_{\text{max}}]_{k}\geq[\mathbf{u}_{\text{max}}]_{k}$ and $\mathbf{v}_{\text{max}}]_{k}=[\mathbf{v}_{\pi(i)}]_{k}$, we have that:
$\big([\mathbf{v}_{\text{max}}]_{k}-[\mathbf{u}_{\text{max}}]_{k}\big)^{2}$ $=\big([\mathbf{v}_{\pi^{*}(i)}]_{k}-[\mathbf{u}_{\text{max}}]_{k}\big)^{2}$
$=\min_{j\in[M]}\big([\mathbf{v}_{\pi^{*}(i)}]_{k}-[\mathbf{u}_{j}]_{k}\big)^{2}$ (3)
$\leq\big([\mathbf{v}_{\pi^{*}(i)}]_{k}-[\mathbf{u}_{i}]_{k}\big)^{2}$
Likewise, $\forall k\in[d]$ where $\mathbf{u}_{\text{max}}]_{k}>[\mathbf{v}_{\text{max}}]_{k}$ and $\mathbf{u}_{\text{max}}]_{k}=[\mathbf{u}_{i}]_{k}$, we have that:
@@
$\leq\big([\mathbf{v}_{\pi^{*}(i)}]_{k}-[\mathbf{u}_{i}]_{k}\big)^{2}$
Then, for each index pair $(\pi^{*}(i),i)$, let $\mathcal{S}_{\pi^{*}(i)}=\{j\colon[\mathbf{v}_{\text{max}}]_{j}=[\mathbf{v}_{\pi^{*}(i)}]_{j},j\in[d]\}$ and $\mathcal{S}_{i}=\{j\colon[\mathbf{u}_{\text{max}}]_{j}=[\mathbf{u}_{i}]_{j},j\in[d]\}$ denote the sets of coordinates at which $\mathbf{v}_{\pi^{*}(i)}$ and $\mathbf{u}_{i}$, respectively, attain the coordinate-wise maximum over their corresponding multisets. Then, from equation B.4.5, and equation B.4.5, we have that:
$\sqrt{\sum_{k\in\mathcal{S}_{\pi^{*}(i)}}\big([\mathbf{v}_{\text{max}}]_{k}-[\mathbf{u}_{\text{max}}]_{k}\big)^{2}+\sum_{k\in\mathcal{S}_{i}\setminus\mathcal{S}_{\pi^{*}(i)}}\big([\mathbf{v}_{\text{max}}]_{k}-[\mathbf{u}_{\text{max}}]_{k}\big)^{2}}\leq\|\mathbf{v}_{\pi^{*}(i)}-\mathbf{u}_{i}\|$
Based on the above, we have that:
$\Big\|f_{\textsc{max}}(X)-f_{\textsc{max}}(Y)\Big\|$ $=\sqrt{\big([\mathbf{v}_{\text{max}}]_{1}-[\mathbf{u}_{\text{max}}]_{1}\big)^{2}+\big([\mathbf{v}_{\text{max}}]_{2}-[\mathbf{u}_{\text{max}}]_{2}\big)^{2}+\ldots+\big([\mathbf{v}_{\text{max}}]_{d}-[\mathbf{u}_{\text{max}}]_{d}\big)^{2}}$
$\leq\|\mathbf{v}_{\pi(1)}-\mathbf{u}_{1}\|+\|\mathbf{v}_{\pi(2)}-\mathbf{u}_{2}\|+\ldots+\|\mathbf{v}_{\pi(M)}-\mathbf{u}_{M}\|$
@@
B.4.6 The max function is Lipschitz continuous with respect to EMD
Let $\mathcal{X}$ denote a set that contains multisets of vectors of equal cardinalities, i. e., $|X|=M$ and $\in\mathcal{S}_{M}(\mathbb{R}^{d}), X𝒳\forall$ $\in\mathcal{X}$ where $M\in\mathbb{N}$. Let $X,Y\in\mathcal{X}$ denote two multisets. We have shown in subsection B.4.5 above that:
$\Big\|f_{\textsc{max}}(X)-f_{\textsc{max}}(Y)\Big\|\leq d_{M}(X,Y)$
@@ B.4.6 The max function is Lipschitz continuous with respect to EMD
B.5 Proof of Proposition 3.3
B.5.1 The att mechanism is not Lipschitz continuous with respect to EMD
Suppose that the att mechanism is Lipschitz continuous with respect to EMD. Let $L>0$ be given. Let also $\epsilon>0$ and $c>\frac{2(1+\exp(d\epsilon))L\epsilon}{\exp(d\epsilon)-1}$. Let $X=\{\mathbf{v}_{1},\mathbf{v}_{2}\}$, $Y=\{\mathbf{u}_{1},\mathbf{u}_{2}\}$ be two multisets, each consisting of $2$ vectors of dimension $d$. Then, suppose that $\mathbf{v}_{1}=\mathbf{u}_{1}=(c,c,\ldots,c)^{\top}$, $\mathbf{v}_{2}=(\epsilon,\epsilon,\ldots,\epsilon)^{\top}$ and $\mathbf{u}_{2}=(-\epsilon,-\epsilon,\ldots,-\epsilon)^{\top}$. Clearly, we have that $d_{EMD}(X,Y)=\frac{1}{2}\|\mathbf{v}_{2}-\mathbf{u}_{2}\|=\sqrt{d}\epsilon$. Let also $\mathbf{W}=-\mathbf{I}$ and $\mathbf{q}=\mathbf{1}$ where $\mathbf{I}$ denotes the $d\times d$ identity matrix and $\mathbf{1}$ denotes the $d$-dimensional vector of all ones. We define $g$ as the ReLU function. The attention coefficients of the elements of $X$ are equal to:
$\alpha_{1}^{X}$ $=\frac{\exp\big(\mathbf{1}^{\top}\textsc{ReLU}(-\mathbf{I}\,\mathbf{v}_{1})\big)}{\sum_{j=1}^{2}\exp\big(\mathbf{1}^{\top}\textsc{ReLU}(-\mathbf{I}\,\mathbf{v}_{j})\big)}=\frac{\exp(0)}{\exp(0)+\exp(0)}=\frac{1}{2}$
$\alpha_{2}^{X}$ $=\frac{\exp\big(\mathbf{1}^{\top}\textsc{ReLU}(-\mathbf{I}\,\mathbf{v}_{2})\big)}{\sum_{j=1}^{2}\exp\big(\mathbf{1}^{\top}\textsc{ReLU}(-\mathbf{I}\,\mathbf{v}_{j})\big)}=\frac{\exp(0)}{\exp(0)+\exp(0)}=\frac{1}{2}$
The attention coefficients of the elements of $Y$ are equal to:
$\alpha_{1}^{Y}$ $=\frac{\exp\big(\mathbf{1}^{\top}\textsc{ReLU}(-\mathbf{I}\,\mathbf{u}_{1})\big)}{\sum_{j=1}^{2}\exp\big(\mathbf{1}^{\top}\textsc{ReLU}(-\mathbf{I}\,\mathbf{u}_{j})\big)}=\frac{\exp(0)}{\exp(0)+\exp(d\epsilon)}=\frac{1}{1+\exp(d\epsilon)}$
$\alpha_{2}^{Y}$ $=\frac{\exp\big(\mathbf{1}^{\top}\textsc{ReLU}(-\mathbf{I}\,\mathbf{u}_{2})\big)}{\sum_{j=1}^{2}\exp\big(\mathbf{1}^{\top}\textsc{ReLU}(-\mathbf{I}\,\mathbf{u}_{j})\big)}=\frac{\exp(d\epsilon)}{\exp(0)+\exp(d\epsilon)}=\frac{\exp(d\epsilon)}{1+\exp(d\epsilon)}$
We then have that:
$\Big\|f_{\textsc{att}}(X)-f_{\textsc{att}}(Y)\Big\|$ $=\bigg\|\sum_{i=1}^{2}\alpha_{i}^{X}\mathbf{v}_{i}-\sum_{j=1}^{2}\alpha_{j}^{Y}\mathbf{u}_{j}\bigg\|$
$=\bigg\|\frac{1}{2}\mathbf{v}_{1}+\frac{1}{2}\mathbf{v}_{2}-\frac{1}{1+\exp(d\epsilon)}\mathbf{u}_{1}-\frac{\exp(d\epsilon)}{1+\exp(d\epsilon)}\mathbf{u}_{2}\bigg\|$
$=\bigg\|\frac{\exp(d\epsilon)-1}{2(1+\exp(d\epsilon))}(c,c,\ldots,c)^{\top}+\frac{3\exp(d\epsilon)+1}{2(1+\exp(d\epsilon))}(\epsilon,\epsilon,\ldots,\epsilon)^{\top}\bigg\|$
$>\bigg\|\frac{\exp(d\epsilon)-1}{2(1+\exp(d\epsilon))}(c,c,\ldots,c)^{\top}\bigg\|$
$>\bigg\|\frac{\exp(d\epsilon)-1}{2(1+\exp(d\epsilon))}\bigg(\frac{2(1+\exp(d\epsilon))L\epsilon}{\exp(d\epsilon)-1},\frac{2(1+\exp(d\epsilon))L\epsilon}{\exp(d\epsilon)-1},\ldots,\frac{2(1+\exp(d\epsilon))L\epsilon}{\exp(d\epsilon)-1}\bigg)^{\top}\bigg\|$
$=\bigg\|L(\epsilon,\epsilon,\ldots,\epsilon)^{\top}\bigg\|$
$=L\sqrt{\underbrace{\epsilon^{2}+\epsilon^{2}+\ldots+\epsilon^{2}}_{d\text{ times}}}$
$=L\sqrt{d}\epsilon$
$=L\,d_{EMD}(X,Y)$